OpenStudy (anonymous):

What is the y coordinate of the y−intercept of the line that passes through the points (−2, −2) and (2, 4)?

OpenStudy (anonymous):

First take out slope m= (2+2)/(4+2)=4/6=2/3 use equation (y-y1)=m(x-x1) taking x1=-2 and y1=-2 y+2=2(x+2)/3 3(y+2)=2x+4 3y+6=2x+4 3y=2x+4-6 3y=2x-2 y=(2x/3)-(2/3) this equation is in slope intercept form y=mx+c so c=-2/3 is your answer